# Towards Human-Level 3D Relative Pose Estimation: Generalizable, Training-Free, with Single Reference

### v1

@@ VI Discussions and Conclusions
Limitations and Future Works. Our method has the following two limitations. Firstly, our method necessitates the depth information of the reference object as an input. Although this is a one-time requirement per object, the need for depth data can restrict the applicability of our method where a depth sensor is absent. To acquire the depth of the reference image, we evaluated several advanced monocular depth estimation algorithms, including midas ; bhat2023zoedepth ; dpt ; depthanything . However, we found that these methods often struggle to generalize across different object types. Despite this, our empirical results, presented in Table S1 of the supplementary materials, demonstrate that our method remains robust with imprecise depth (simulated by adding noise to the ground-truth depth). This suggests that the current limitations are likely to be overcome once an object-generalizable depth estimator becomes available.
@@
Conclusions. In this paper, we addressed the challenging generalizable relative pose estimation under a rigorous circumstance with only a single RGB-D reference and single RGB query pair as input, and the pose label is not a priori. We establish our label- and training-free method following the render-and-compare paradigm, by exploiting 1) the 2.5D (i.e., RGB-D) rotatable reference mesh, 2) the semantic maps of both query and reference (extracted by a pretrained large vision model DINOv2), and 3) a differentiable renderer to produce and back-propagate losses to refine the relative pose. We carried out extensive experiments on the LineMOD, LM-O, and YCB-V datasets. The results demonstrate that our label/training-free approach surpasses the performance of state-of-the-art supervised methods, particularly excelling under the rigorous Acc@5/10/15 metrics.

### v2

@@ VI Discussions and Conclusions
Limitations and Future Works. Our method has the following two limitations. Firstly, our method necessitates the depth information of the reference object as an input. To acquire the metric depth of the reference image, we evaluated a state-of-the-art monocular depth estimation algorithm yang2024depth . The results and discussions in Sect. S3 of the supplementary shows that the metric depth estimation occasionally fails to generalize across different datasets, primarily stemming from the inherent metric scale ambiguities under varying camera parameters and diverse objects. This suggests that the current limitations are likely to be overcome once a generalizable metric depth estimator becomes available. Despite this, we note that depth sensors are commonly used in our primary application domain, i.e., robotics. Our empirical results, presented in Table S1 of the supplementary materials, demonstrate that our method remains robust with imprecise depth obtained by a noisy depth sensor (simulated by adding noise to the ground-truth depth).
@@
Conclusions. In this paper, we addressed the challenging generalizable relative pose estimation under a rigorous circumstance with only a single RGB-D reference and single RGB query pair as input, and the pose label is not a priori. We establish our label- and training-free method following the render-and-compare paradigm, by exploiting i) the 2.5D (i.e., RGB-D) rotatable reference mesh, ii) the semantic maps of both query and reference (extracted by a pretrained large vision model DINOv2), and iii) a differentiable renderer to produce and back-propagate losses to refine the relative pose. We carried out extensive experiments on the LineMOD, LM-O, and YCB-V datasets. The results demonstrate that our label/training-free approach surpasses the performance of state-of-the-art supervised methods, particularly excelling under the rigorous Acc@5/10/15 metrics.